Joe invested a portion of $6000 at 4.5% interest and the balance at 6% interest. How much did he invest at 6% if his total income from both investments is $279?
Answer by checkley77(12569)

.06X+.045(6,000-X)=279
.06X+270-.045X=279
.015X=279-270
.015X=9
X=9/.015
X=$600 WAS INVESTED @ 6%.
6,000-600=5,400 INVESTED @ 4.5%
PROOF:
.06*600+.045*5,400=279
36+243=279
279=279